OpenStudy (anonymous):

Solve for the variable. 5 - x = 16

OpenStudy (skullpatrol):

{-11}

OpenStudy (skullpatrol):

5 - x = 16 (-5) + 5 - x = (-5) + 16 0 - x = 11 -x = 11 (-1)(-x) = (-1)(11) x = -11 {-11}